how many protons must flow through ATP synthase per day in a resting human being requiring 85 kg of ATP per day?
Answer #1

3 protons are require for the production of 1 ATP molecule

molar mass of ATP = 507 g/mol

so 85 Kg of ATP = 85000/507 = 167.65 moles of ATP

167.65 moles of ATP contains = 167.65*6.023*10^23 = 1.01*10^26 ATP molecules

1 molecule of ATP require 3 Protons

so number of Protons required = 3.03*10^26 number of protons